A 46 kg sample of water absorbs 354 kj of heat. If the water was initially at 21.1 C, what is its final temperature?
1
Expert's answer
2012-10-22T11:53:50-0400
q = mc delta t
m = 46 000 g
c = 4.186
q = 354 000 j

delta t = q/mc = 354000/(46000*4.186) = 1.84

finel t = 21.1 + 1.84 = 22.94 C